OpenStudy (anonymous):

Solve using substitution: 5y − x = -2 -2x = 6 − 10y A. No solution B. Infinitely many solutions C. (3, -1) D. (-8, -2)

OpenStudy (zzr0ck3r):

same thing as the last do you not see how I did it?

OpenStudy (anonymous):

yes but still dont understand. sorry /:

OpenStudy (zzr0ck3r):

x = 2 + 5y -2(2+5y) = 6 - 10y -4 -10y = 6 - 10y no solutions because -4 != 6

OpenStudy (anonymous):

was that wrong which i did???

OpenStudy (anonymous):

i dont know yet its a review packet i have to do @sradha

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

let y = 0 for the first one what does x = 2 use -4 = 6 − 10y

OpenStudy (anonymous):

now what does y =?

OpenStudy (anonymous):

10y = 10

OpenStudy (zzr0ck3r):

no, there are no solutions. And that is not substitution

OpenStudy (anonymous):

ahhh oppsy

OpenStudy (anonymous):

well i tried

OpenStudy (zzr0ck3r):

I took the first equation and changed it around to be x = somehting( you could have made it y = something) then pluf that in for x (or y ) in the second equation. the algebra works out to be -4 = 6 thus there are no solutions

OpenStudy (anonymous):

theres no solution for it............

OpenStudy (zzr0ck3r):

you can always graph these and see where the intersect (these two dont)